# Evaluating Long Range Dependency Handling in Code Generation Models using Multi-Step Key Retrieval

### v1

@@ Appendix A Detailed Results: Experiment 1
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 27.0 18.5 6.4 12.8 6.2 1.2 1.1 0.6 0.4 StarCoder2-7B 54.3 36.2 13.4 26.5 23.0 5.4 8.6 3.9 1.1 StarCoderBase-1B 2.9 1.5 0.6 1.1 0.5 0.2 0.3 0.1 0.0 StarCoderBase-7B 37.5 19.2 9.1 26.9 11.7 1.6 16.1 7.3 0.5 StarCoderBase-15.5B 50.5 36.6 16.1 35.6 21.9 7.3 27.5 13.3 4.6
Table 6: Overall model performance vs. number of forward references. 5 distractors. Note that forward references cannot exist in one-step tasks.
Two-step Task
Context Size 2k 4k 8k # Forward References 0 1 0 1 0 1 Mistral-7B-v0.1 60.1 21.1 46.4 22.5 8.0 9.2 StarCoder2-7B 73.5 57.6 69.4 56.5 28.1 26.8 StarCoderBase-1B 55.3 39.4 49.7 33.9 65.6 39.3 StarCoderBase-7B 90.0 38.1 87.2 34.1 76.0 37.7 StarCoderBase-15.5B 85.5 63.4 82.3 49.2 63.4 28.3
Three-step Task
Context Size 2k 4k 8k # Forward References 0 1 2 0 1 2 0 1 2 Mistral-7B-v0.1 28.1 24.4 16.8 19.8 22.3 16.9 2.7 11.7 10.3 StarCoder2-7B 61.5 60.5 45.7 46.2 48.1 37.5 14.4 23.1 16.0 StarCoderBase-1B 32.3 29.5 18.1 29.5 28.5 14.9 19.8 23.8 8.2 StarCoderBase-7B 63.4 41.3 18.8 65.4 42.0 20.5 41.9 35.6 15.9 StarCoderBase-15.5B 57.2 50.6 37.2 54.0 39.7 24.9 34.5 21.0 10.5
Concatenation Task
Context Size 2k 4k 8k # Forward References 0 1 2 0 1 2 0 1 2 Mistral-7B-v0.1 11.8 4.3 3.1 2.4 0.6 0.6 0.8 0.2 0.2 StarCoder2-7B 17.7 14.7 7.9 6.4 5.4 4.2 1.8 0.7 0.7 StarCoderBase-1B 1.6 0.3 0.0 0.5 0.1 0.0 0.0 0.0 0.0 StarCoderBase-7B 13.9 9.3 4.0 2.0 1.6 1.2 0.6 0.6 0.3 StarCoderBase-15.5B 22.9 14.2 11.0 10.1 6.1 5.8 6.8 4.5 2.6
@@
Appendix B Detailed Results: Call Graph Comment Experiment
We include details of model performance (accuracy@3 percent scores) for the full-sentence template variant of the call graph experiment.
Table 7: Call graph comment performance by model.
@@ Appendix B Detailed Results: Call Graph Comment Experiment
Figure 14: Effect of call graph comments on Mistral-7B-v0.1.
Figure 15: Effect of call graph comments on Starcoder2-7B performance.
Figure 16: Effect of call graph comments on StarCoderBase-1B performance.
Figure 17: Effect of call graph comments on StarCoderBase-7B performance.
Figure 18: Effect of call graph comments on StarCoderBase-15.5B performance.
Appendix C Hyperparameters for Generation
@@
Appendix D Confidence Intervals
@@
Appendix E Compute Information
We ran all experiments on machines with a single A100 GPU with 80GB of VRAM on a cloud provider. Experiments were run over the course of about a month and we estimate ~1000 GPU hours were used including runs that had bugs in them and additional experiments that are not part of the paper. We use FlashAttention [6] to improve memory usage and latency of generation.
Appendix F Prompt Samples for Multi Step Key Retrieval
@@
Listing 1: Sample prompt for two-step retrieval task with 2000 token context size and 5 distractors.
Appendix G Prompt Samples for Call Graph Comments Experiment
@@
Listing 2: Full sentence call graph comments for three step retrieval. 5 distractors. HumanEval functions excluded for brevity.
@@
Listing 3: Function names only call graph comments for three step retrieval. 5 distractors. HumanEval functions excluded for brevity.

### v2

@@ Appendix B Detailed Results: Experiment 1
Context Size 2k 4k 8k # Distractors 0 1 5 0 1 5 0 1 5 Mistral-7B-v0.1 27.0 18.5 6.4 12.8 6.2 1.2 1.1 0.6 0.4 StarCoder2-7B 54.3 36.2 13.4 26.5 23.0 5.4 8.6 3.9 1.1 StarCoderBase-1B 2.9 1.5 0.6 1.1 0.5 0.2 0.3 0.1 0.0 StarCoderBase-7B 37.5 19.2 9.1 26.9 11.7 1.6 16.1 7.3 0.5 StarCoderBase-15.5B 50.5 36.6 16.1 35.6 21.9 7.3 27.5 13.3 4.6 DeepSeekCoder-6.7B-base 60.1 50.1 28.4 54.2 33.3 10.0 47.4 20.4 5.8 GPT-4o-mini 85.4 78.8 71.5 74.4 84.0 54.4 73.0 62.2 36.4 GPT-4o 97.6 98.7 97.5 93.2 97.3 96.1 94.8 92.4 91.4
@@
StarCoder2-7B
DeepSeekCoder 6.7B-base
GPT-4o-mini
GPT-4o
Appendix C Incorrect Response Analysis
Here we present analyses of incorrect responses broken down by model. We also present details of the edit distance analysis.
Table 7: Percentage of distractor answers across all experiments with distractors
Prompt Size 2k 4k 8k Mean
Mistral-7B-v0.1 10.5 9.4 5.3 8.4
StarCoder2-7B 16.7 19.4 12.9 16.3
StarCoderBase-1B 11.5 11.0 14.2 12.2
StarCoderBase-7B 7.7 9.5 11.7 9.6
StarCoderBase-15.5B 11.5 11.3 10.1 11.0
DeepSeekCoder-6.7B-base 16.0 17.5 15.1 16.2
GPT-4o-mini 5.2 7.3 11.1 7.9
GPT-4o 0.9 1.9 2.7 1.8
Table 8: Percentage of incorrect answers that are distractor values
Prompt Size 2k 4k 8k Mean
Mistral-7B-v0.1 14.3 12.5 6.0 10.9
StarCoder2-7B 30.8 35.0 16.9 27.6
StarCoderBase-1B 15.9 16.2 22.2 18.1
StarCoderBase-7B 12.5 15.6 20.8 16.3
StarCoderBase-15.5B 19.8 17.9 15.5 17.7
DeepSeekCoder-6.7B-base 36.4 34.6 29.5 33.5
GPT-4o-mini 31.0 33.8 35.5 33.4
GPT-4o 25.8 29.1 24.2 26.3
Table 9: Percentage of incorrect responses in the concatenation task where the response is one of the two strings that should have been concatenated.
Prompt Size 2k 4k 8k Mean
Mistral-7B-v0.1 20.7 26.7 11.5 19.6
StarCoder2-7B 21.8 33.5 30.0 28.5
StarCoderBase-1B 44.9 43.0 44.2 44.0
StarCoderBase-7B 40.9 50.9 51.6 47.8
StarCoderBase-15.5B 30.3 33.1 32.5 32.0
DeepSeekCoder-6.7B-base 23.4 30.4 34.7 29.5
GPT-4o-mini 28.1 26.7 32.3 29.0
GPT-4o 13.7 7.6 8.1 9.8
Table 10: Levenshtein Distance for incorrect responses.
Prompt Size 2k 4k 8k Mean
Mistral-7B-v0.1 10.8 12.3 15.1 12.7
StarCoder2-7B 10.5 10.5 11.4 10.8
StarCoderBase-1B 9.6 9.4 9.7 9.6
StarCoderBase-7B 9.5 9.7 9.7 9.6
StarCoderBase-15.5B 11.1 11.9 12.3 11.8
DeepSeekCoder-6.7B-base 9.3 9.9 10.2 9.8
GPT-4o-mini 8.5 10.0 12.1 10.2
GPT-4o 9.5 11.4 12.4 11.1
Appendix D Detailed Results: Call Graph Comment Experiment
We include details of model performance (accuracy@3 percent scores) on the call graph experiment.
Table 11: Call graph comment performance by model for full-sentence template.
@@ Appendix D Detailed Results: Call Graph Comment Experiment
DeepSeekCoder-6.7B-base 14.7 17.9 27.2 35.8
GPT-4o-mini 54.1 65.1 71.1 73.3
GPT-4o 95.0 95.3 96.1 94.0
Table 12: Call graph comment performance by model for names only template.
Task Model Comment Type
None Calls Called-By Both
three-step Mistral-7B-v0.1 18.2 15.0 36.1 45.4
StarCoder2-7B 41.6 47.5 72.0 71.8
StarCoderBase-1B 25.0 28.9 51.5 48.6
StarCoderBase-7B 39.0 35.5 80.1 71.6
StarCoderBase-15.5B 36.9 35.5 80.8 63.0
DeepSeekCoder-6.7B-base 58.5 60.0 83.6 80.5
GPT-4o-mini 64.8 66.6 81.4 78.5
GPT-4o 91.6 93.1 97.9 94.9
concatenation Mistral-7B-v0.1 2.7 2.3 4.0 4.1
StarCoder2-7B 6.6 8.5 15.4 11.3
StarCoderBase-1B 0.3 0.2 0.4 0.3
StarCoderBase-7B 3.7 6.9 4.0 11.4
StarCoderBase-15.5B 9.3 11.2 24.6 19.2
DeepSeekCoder-6.7B-base 14.7 15.2 23.4 23.8
GPT-4o-mini 54.1 58.2 68.9 70.5
GPT-4o 95.0 94.6 94.9 95.7
@@
DeepSeekCoder-6.7B-base 0 54.8 84.5 87.9
1 61.7 86.1 91.0
2 49.6 66.7 77.7
GPT-4o-mini 0 53.6 70.0 75.2
1 68.4 84.3 88.1
2 61.4 78.7 84.7
GPT-4o 0 89.8 98.1 97.8
1 93.0 98.3 97.9
2 87.6 96.9 97.7
Figure 15: Effect of call graph comments for three-step task.
Figure 16: Effect of call graph comments for concatenation task.
Appendix E Hyperparameters for Generation
@@
Appendix F Confidence Intervals
@@
Appendix G Compute Information
We ran all experiments on machines with a single A100 GPU with 80GB of VRAM on a cloud provider. Experiments were run over the course of about a month and we estimate ~1000 GPU hours were used including runs that had bugs in them and additional experiments that are not part of the paper. We use FlashAttention Dao et al. (2022) to improve memory usage and latency of generation.
Appendix H Prompt Samples for Multi Step Key Retrieval
@@
Listing 1: Sample prompt for two-step retrieval task with 2000 token context size and 5 distractors. rdcxoi_135343 is the key function and egllun_467846 is the value function.
Appendix I Prompt Samples for Call Graph Comments Experiment
@@
assert wwzfoa_904885() ==
Listing 2: Full sentence call graph comments for three step retrieval. 5 distractors. HumanEval functions excluded for brevity. wwzfoa_904885 is the key function, bweckw_860527_nykiyp and gjobme_651008_tymmij are value functions
@@
Listing 3: Function names only call graph comments for three step retrieval. 5 distractors. HumanEval functions excluded for brevity. wwzfoa_904885 is the key function, bweckw_860527_nykiyp and gjobme_651008_tymmij are value functions